OpenStudy (anonymous):

7^3x-17=2,401. solve and express each side as a power of the same base and then equating exponents

OpenStudy (anonymous):

\[7^{3x+17}=2,401\]

OpenStudy (anonymous):

what's 7^2, 7^3, 7^4

OpenStudy (anonymous):

7^2=49 7^3=343 7^4=2401 So we may rewrite 2401 as 7^4

OpenStudy (anonymous):

\[7^{3x+17}=7^4\]